# Towards Aligned Canonical Correlation Analysis: Preliminary Formulation and Proof-of-Concept Results

### 01 Dec 2023 v1

@@ Abstract.
Canonical Correlation Analysis (CCA) has been widely applied to jointly embed multiple views of data in a maximally correlated latent space. However, the alignment between various data perspectives, which is required by traditional approaches, is unclear in many practical cases. In this work we propose a new framework Aligned Canonical Correlation Analysis (ACCA), to address this challenge by iteratively solving the alignment and multi-view embedding.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: 19th International Workshop on Mining and Learning with Graphs; August $X$, 2023; Long Beach, CA
@@ 1. Introduction
Canonical Correlation Analysis (Harold, 1936; Kettenring, 1971) is a classical model which, given two different views of the same set of entities, e.g., two different bipartite graphs of (user, product) and (user, video) interactions or different feature representations for those entities in general, seeks to project those entities (users) in a low-dimensional space where the different projected views are maximally correlated.
In traditional CCA-style analysis, we assume that entities across multiple views have one-to-one correspondence , and there is a wealth of algorithms that study different formulations for solving the problem of projecting those views in that desired maximally correlated space, both linearly and non-linearly (Andrew et al., 2013). However, in many real-world applications, the data of different views are generated by different resources, respectively, which potentially causes the imperfect alignment of multiple views corresponding to the same entity, e.g, the multiple medical information always be recorded by different hospitals correspondingly that the alignment of those records from the same patient is unknown. In such cases, to preserve the performance of the CCA analysis, a proper estimation of multi-view alignment is necessary. To address this problem, inspired by recent work (Wu et al., 2022) in the related problem of misaligned joint tensor factorization, we proposed a new formulation, Aligned Canonical Correlation Analysis (ACCA), which seeks to iteratively identify the best entity alignment and latent embedding for multiple views of one dataset. We derive an Alternating Optimization algorithm and present preliminary results to demonstrate the feasibility of our framework
@@ 2. Background
[…]
@@ 3.1. Proposed Formulation for ACCA
Consider two centered datasets $\mathbf{X}\in\mathbb{R}^{D_{x}\times N}$ and $\mathbf{Y}\in\mathbb{R}^{Dy\times N}$, and assume the entity alignment between the columns of the two datasets is unknown $\bar{\mathbf{P}}\in\mathbf{R}^{N\times N}$, our goal is to learn the latent component representation $\mathbf{S}$ and predict the alignment matrix $\bar{\mathbf{P}}$ simultaneously. Let’s denote the estimation of $\bar{\mathbf{P}}$ to be ${\mathbf{P}}\in\mathbf{R}^{N\times N}$. Ideally, $\mathbf{P}$ should be a permutation matrix satisfying: (1) $\mathbf{P}$ is a binary matrix; (2) the sum of each row is one; and (3) the sum of each column is one, where all the constraints can be used to prove that $\mathbf{P}$ is an orthogonal matrix. Mathematically, we will minimize $\left\|\mathbf{U}{\mathbf{X}}-\mathbf{S}\right\|_{F}^{2}+\left\|\mathbf{V}{\mathbf{Y}}\mathbf{P}-\mathbf{S}\right\|_{F}^{2}$ under the above three constraints as well as the constraint from CCA, i.e., $\mathbf{SS}^{\top}=\mathbf{I}$. However, solving such an optimization problem is challenging. Motivated by earlier work (Wu et al., 2022), instead of directly solving for a permutation matrix, which is computationally prohibitive, we define a number of constraints which describe different aspects of a permutation matrix without strictly enforcing it to be one, allow us to have a tractable optimization solution. We, thus, relax the constraints on $\mathbf{P}$ leading to our proposed ACCA model:
@@
(3) S. T. $\mathbf{SS}^{\top}=\mathbf{I},\text{(uncorrelatedness)}$
(4) $0\leq p_{i,j}\leq 1,\forall i,j,\text{(nonnegativity)}$
(5) $\sum_{j=1}^{N}p_{i,j}=1,\forall i,\text{(row-wise sum )}$
(6) $H(\mathbf{p}_{i})\leq\lambda,\forall i\text{( entropy)}$
where $p_{i,j}$ is the $(i,j)$-th entry of $\mathbf{P}$, $\mathbf{p}_{i}$ is the $i$-th row of $\mathbf{P}$, $H(\mathbf{p}_{i})$ is the entropy of $\mathbf{p}_{i}$ by viewing the $N$ entries of $\mathbf{p}_{i}$ as a discrete probability distribution, and the hyperparamters $\gamma_{1}$, $\gamma_{2}$, and $\lambda$ are nonnegative. It’s worth to mention that enforcing the low entropy of $\mathbf{p}_{i}$ guarantees that the distribution is far away from uniform distribution and close to a deterministic distribution as an ideal $\mathbf{p}_{i}$ has a single $1$ value with the rest to be 0s. Furthermore, the second and third terms in the objective function in Eq. (2) promote the orthogonality of $\mathbf{P}$ since an ideal permutation matrix satisfies that the columns/rows are orthonormal to each each.
@@ 3.2. Alternating Optimization for ACCA
To solve the ACCA formulation, we will adopt alternating optimization method. Specifically, we will iteratively seek for the optimal CCA related variables ($\mathbf{U}$, $\mathbf{V}$, and $\mathbf{S}$) while fixing $\mathbf{P}$ to be the update from the previous iteration, and vice versa. When looking for CCA related variables ($\mathbf{U}$, $\mathbf{V}$, and $\mathbf{S}$), the sub-optimization problem is reduced to be the traditional CCA formulation by substituting $\mathbf{Y}$ in (1) with $\mathbf{Y}\mathbf{P}$. In the subproblem of optimizing $\mathbf{P}$, we use the scipy.optimize.minimize solver in Python  [^1] https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.minimize.html . The proposed framework is summarized in Algorithm 1.
@@ 4.1. Synthetic Data Generation
We first generate the groundtruth latent representation of the two datasets, namely $\mathbf{Z}\in\mathbb{R}^{\bar{d}\times N}$, where the columns of $\mathbf{Z}$ are $N$ i.i.d. samples drawn from multivariate normal distribution with zero mean and identity covariance of size $\bar{d}\times\bar{d}$. Next, two aligned datasets $\mathbf{X}$ and $\bar{\mathbf{Y}}\in\mathbb{R}^{D_{y}\times N}$ are generated from their shared latent representation $\mathbf{Z}$ through two independent random projections: $\mathbf{X}=\mathbf{W}\mathbf{Z}$ and $\mathbf{Y}=\mathbf{Q}\mathbf{Z}$ where $\mathbf{W}\in\mathbb{R}^{D_{x}\times\bar{d}}$ and $\mathbf{Q}\in\mathbb{R}^{D_{y}\times\bar{d}}$. For each experiment, the groundtruth $\bar{\mathbf{P}}$ is a random permutation matrix with only one entry in each row and column to be $1$ and the rest to be 0s. Next, we have two unaligned datasets: $\mathbf{X}$ and $\mathbf{Y}=\bar{\mathbf{Y}}\bar{\mathbf{P}}$. The involved parameters are set as follows: $N=20$, $\bar{d}=2$, $d=7$, $D_{x}=15$, and $D_{y}=10$.
@@ 4.2. Experimental Results
Next, we visualize the alignment performance with respect to different values of the hyperparameter $\lambda$ in Figure 3 where we plot the real permutation matrix $\bar{\mathbf{P}}$ and the estimated $\mathbf{P}$ as gray-scale images with darker grid blocks representing higher values of the corresponding entries of $\bar{\mathbf{P}}$ or $\mathbf{P}$. As uniform distribution leads to the highest entropy, $\lambda$ can not exceed $log(N)$ ($N\times 1/N\times log(1/N)$). With $\lambda$ increasing, more nonzero entries are showing up in $\mathbf{P}$ as expected. With proper setup of entropy bound hyperparameter, the performance of ACCA will be further improved, with the comparison of prediction accuracies related to different entropy cases in Figure 3.
@@ Acknowledgements.
The authors would like to thank Yunshu Wu for initial discussions. Research was supported by the National Science under CAREER grant no. IIS 2046086 and CREST Center for Multidisciplinary Research Excellence in Cyber-Physical Infrastructure Systems (MECIS) grant no. 2112650 and by the US Department of Transportation under University Transportation Center (UTC) on Railway Safety. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding parties. $\sum_{i=1}^{N}P_{ij}=1$, for $\forall j\in[N]$

### 08 Dec 2023 v2

@@ Abstract.
Canonical Correlation Analysis (CCA) has been a tried and tested analytical model which seeks to jointly embed two or more views of the data in a maximally correlated latent space. In addition to a powerful data mining tool, CCA has many implications for cutting-edge self-supervised representation learning approaches, as one can cast a number of recent approaches as variants of that model. In this work we explore what happens when a fundamental assumption of that model breaks: what if the alignment across different views of the data is unknown? Typically, we would first attempt to align the different views, and subsequently apply CCA or any other model in order to embed the data views. Can we do better if we align and embed at the same time? In this work, we seek to jointly solve the alignment and embedding by formulating and solving both problems under the same umbrella of Aligned Canonical Correlation Analysis (ACCA). We present a preliminary formulation and alternating optimization algorithm and proof-of-concept results.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Canonical Correlation Analysis (CCA) (Harold, 1936; Kettenring, 1971) is a classical model which, given two different views of the same set of entities, e.g., two different bipartite graphs of (user, product) and (user, video) interactions or different feature representations for those entities in general, seeks to project those entities (users) in a low-dimensional space where the different projected views are maximally correlated. Essentially, CCA can jointly embed heterogeneous datasets in a common low-dimensional space, as it can be extended to more than two views (Chen et al., 2019, 2022). Even though CCA has been in and out of the spotlight for many decades and has been around for quite a long time, it is still extremely relevant, not only as a standalone data mining tool, but also due to the fact that cutting-edge self-supervised representation learning techniques, such as Barlow Twins (Zbontar et al., 2021; Bielak et al., 2022; Zhang et al., 2021; Balestriero et al., 2023), can be essentially seen as variations of CCA, where the goal is to embed two views of the data (the original view and the augmentation) in a latent space where correlation is maximized.
Traditionally, in CCA-style analysis, we assume that entities across views have one-to-one correspondence across the two or more views of the data, and there is a wealth of algorithms that study different formulations for solving the problem of projecting those views in that desired maximally correlated space, both linearly and non-linearly (Andrew et al., 2013). What if, however, this one-to-one correspondence is unknown? In this case, we are faced with two problems: (1) entity alignment and (2) CCA embedding. Motivated by recent results (Wu et al., 2022) in the related problem of misaligned joint tensor factorization, it turns out that formulating and solving the alignment and embedding problems jointly yields better results than solving each problem separately in multiple steps, as it appears that the two sub-problems work synergistically to produce better quality alignment and embeddings. In this work, we explore this type of formulation for CCA and we propose a new formulation, the Aligned Canonical Correlation Analysis (ACCA), where we seek to jointly compute the alignment and the embedding space.
@@ 2. Background
(1) ${\textbf{min}}_{\mathbf{U,V,S}}||\mathbf{UX}-\mathbf{S}||^{2}_{F}+||\mathbf{VY}-\mathbf{S}||^{2}_{F}$
@@ 3.1. Proposed Formulation for ACCA
Consider two centered datasets $\mathbf{X}\in\mathbb{R}^{D_{x}\times N}$ and $\mathbf{Y}\in\mathbb{R}^{Dy\times N}$ are two views in one dataset, and the alignment between the columns of the two views, denoted as $\bar{\mathbf{P}}\in\mathbf{R}^{N\times N}$, is unknown. Our goal is to learn the latent component representation $\mathbf{S}$ and predict the alignment matrix $\bar{\mathbf{P}}$ iteratively. Let’s denote the estimation of $\bar{\mathbf{P}}$ as ${\mathbf{P}}\in\mathbf{R}^{N\times N}$. Theoretically, $\mathbf{P}$ should be a (binary) permutation matrix, and the sum of row or column is one, which shows that $\mathbf{P}$ is an orthogonal matrix. Mathematically, we will minimize $\left\|\mathbf{U}{\mathbf{X}}-\mathbf{S}\right\|_{F}^{2}+\left\|\mathbf{V}{\mathbf{Y}}\mathbf{P}-\mathbf{S}\right\|_{F}^{2}$ under the constraints of $\mathbf{P}$ as well as the constraints from CCA, i.e., $\mathbf{SS}^{\top}=\mathbf{I}$. To address the computational limitation in such optimization problem, we define a list of constraints to describe different aspects of a permutation matrix instead of enforcing it to be one, for a tractable optimization solution. By relaxing the constraints on $\mathbf{P}$, the optimization formulation of our proposed ACCA is shown as:
@@
[…]
(6) $H(\mathbf{p}_{i})\leq\lambda,\forall i\text{( entropy)}$)
where $p_{i,j}$ is the $(i,j)$-th entry of $\mathbf{P}$, $\mathbf{p}_{i}$ is the $i$-th row of $\mathbf{P}$, $H(\mathbf{p}_{i})$ is the entropy of $\mathbf{p}_{i}$ by viewing the $N$ entries of $\mathbf{p}_{i}$ as a discrete probability distribution, and the hyperparamters $\gamma_{1}$, $\gamma_{2}$, and $\lambda$ are nonnegative. Enforcing the low entropy of $\mathbf{p}_{i}$ guarantees that the distribution is closed to a deterministic distribution, and the second and third terms in Eq. (2) will promote the orthogonality of $\mathbf{P}$
@@ 3.2. Alternating Optimization for ACCA
To solve the ACCA formulation, as the solution of CCA is dependent on the estimation of permutation matrix, and vice versa, we adopt alternating optimization method, shown in Algorithm 1.
@@ 4.1. Synthetic Data Generation
We first generate the groundtruth latent representation of the two datasets, namely $\mathbf{Z}\in\mathbb{R}^{\bar{d}\times N}$, where the columns of $\mathbf{Z}$ are $N$ i.i.d. samples drawn from multivariate normal distribution with zero mean and identity covariance of size $\bar{d}\times\bar{d}$. Next, two aligned datasets $\mathbf{X}$ and $\bar{\mathbf{Y}}\in\mathbb{R}^{D_{y}\times N}$ are generated from their shared latent representation $\mathbf{Z}$ through two independent random projections: $\mathbf{X}=\mathbf{W}\mathbf{Z}$ and $\mathbf{Y}=\mathbf{Q}\mathbf{Z}$ where $\mathbf{W}\in\mathbb{R}^{D_{x}\times\bar{d}}$ and $\mathbf{Q}\in\mathbb{R}^{D_{y}\times\bar{d}}$. For each experiment, the groundtruth $\bar{\mathbf{P}}$ is a random permutation matrix with only one entry in each row and column to be $1$ and the rest to be 00s. Next, we have two unaligned datasets: $\mathbf{X}$ and $\mathbf{Y}=\bar{\mathbf{Y}}\bar{\mathbf{P}}$. The involved parameters are set as follows: $N=20$, $\bar{d}=2$, $d=7$, $D_{x}=15$, and $D_{y}=10$.
@@ 4.2. Experimental Results
Next, we visualize the alignment performance with respect to different values of the hyperparameter $\lambda$ in Figure 3 where we plot the real permutation matrix $\bar{\mathbf{P}}$ and the estimated $\mathbf{P}$ as gray-scale images with darker grid blocks representing higher values of the corresponding entries of $\bar{\mathbf{P}}$ or $\mathbf{P}$. As uniform distribution leads to the highest entropy, $\lambda$ can not exceed $log(N)$ (=$N\times 1/N\times log(1/N)$). With $\lambda$ increasing, more nonzero entries are showing up in $\mathbf{P}$ as expected. With proper setup of entropy bound hyperparameter, the performance of ACCA will be further improved, with the comparison of prediction accuracies related to different entropy cases in Figure 3.
@@ Acknowledgements.
The authors would like to thank Yunshu Wu for initial discussions. Research was supported by the National Science under CAREER grant no. IIS 2046086 and CREST Center for Multidisciplinary Research Excellence in Cyber-Physical Infrastructure Systems (MECIS) grant no. 2112650 and by the US Department of Transportation under University Transportation Center (UTC) on Railway Safety. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding parties.